# When and How: Learning Identifiable Latent States for Nonstationary Time Series Forecasting

### v1

@@ Abstract
Temporal distribution shifts are ubiquitous in time series data. One of the most popular methods assumes that the temporal distribution shift occurs uniformly to disentangle the stationary and nonstationary dependencies. But this assumption is difficult to meet, as we do not know when the distribution shifts occur. To solve this problem, we propose to learn IDentifiable latEnt stAtes (IDEA) to detect when the distribution shifts occur. Beyond that, we further disentangle the stationary and nonstationary latent states via sufficient observation assumption to learn how the latent states change. Specifically, we formalize the causal process with environment-irrelated stationary and environment-related nonstationary variables. Under mild conditions, we show that latent environments and stationary/nonstationary variables are identifiable. Based on these theories, we devise the IDEA model, which incorporates an autoregressive hidden Markov model to estimate latent environments and modular prior networks to identify latent states. The IDEA model outperforms several latest nonstationary forecasting methods on various benchmark datasets, highlighting its advantages in real-world scenarios.
@@ 1 Introduction
Time series forecasting (Zhou et al., 2021; Lim & Zohren, 2021; Rangapuram et al., 2018; Chatfield, 2000; Zhang, 2003) is known as one of the fundamental problems in machine learning and has achieved pioneering applications in various fields (Bi et al., 2023; Wu et al., 2023; Sezer et al., 2020). However, the inherent non-stationarity of time series data hinders the forecasting models from generalizing on the temporally varying distribution shift.
Several methodologies are proposed to solve this problem, which can be categorized into two types according to the inter-instance and intra-instance temporal distribution shift assumptions. The first type of method assumes that the shift occurs among instances, and each sequence instance is stationary (Li et al., 2023b; Oreshkin et al., 2021). Therefore, instance normalization (Kim et al., 2021) or nonstationary attention mechanism (Liu et al., 2022) are used to remove nonstationary components and compensate for them in prediction. Another type assumes that the environment changes uniformly (Liu et al., 2023c; Surana, 2020). Therefore, some researchers adopt stationarization (Virili & Freisleben, 2000) to remove nonstationarity from time series data. And (Liu et al., 2022) partitions the time-series data into equally-sized and stationary segments and uses the Fast Fourier Transform to select stationary and nonstationary components. In summary, these methods aim to disentangle the stationary and nonstationary dependency. More discussion about related works can be found in Appendix A.
Although current methods mitigate the temporal distribution shift to some extent, the assumptions they require are too strict since each sequence instance or segment may not be stationary when the environment labels are unknown. Figure 1 illustrates toy examples where a nonstationary sine curve is influenced by the nonstationary latent variable (amplitude) and stationary latent variables (frequency and phase). As shown in Figure 1 (a), existing methods split the nonstationary time series into three equal-size subsequences, but the purple and green curves are still nonstationary, making it difficult to disentangle the stationary and nonstationary latent variables. Furthermore, according to Figure 1 (b), even when latent environments are estimated correctly, if the latent variables are entangled, that is, the estimated amplitude and phase are not identified to change simultaneously, the forecasting model can still achieve suboptimal results.
@@
Based on the above examples, an intuitive solution to nonstationary time series forecasting is shown in Figure 1 (c), showing that we should first detect when the temporal distribution shift occurs and then disentangle the stationary/nonstationary states to learn how they change across time. Under this intuition, we find that the latent environments and the stationary/nonstationary variables can be identified theoretically via sufficient observation assumption of nonstationarity. Therefore, we first consider a data generation process for nonstationary time series data. Sequentially, we employ sufficient observations assumption to present an identification theory for latent environments. We further prove that the stationary and nonstationary latent variables are identifiable. Following the theoretical results, we devise a model named IDEA to learn identifiable latent states for nonstationary time series forecasting. Specifically, the proposed IDEA is built on a variational inference framework, incorporating an autoregressive hidden Markov model to estimate latent environments and modular prior network architectures to identify stationary and nonstationary latent variables. Evaluation of simulation and eight real-world benchmark datasets demonstrate the accuracy of latent environment estimation and identification of latent states, as well as the effectiveness of real-world applications.
2 Identifying Distribution of Time Series Data
@@ 2.1 Data Generation Process for Time Series Data
To illustrate how we address the nonstationary time series forecasting problem, we begin with the data generation process as shown in Figure 2. Suppose that we have time series data with discrete time steps, ${\mathbf{x}}=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{T}\}$, where ${\mathbf{x}}_{t}\in\mathbb{R}$ are generated from latent variables ${\mathbf{z}}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{n}$ by an invertible and non-linear mixing function $g$.
$\vspace{-2.5mm}\small{\mathbf{x}}_{t}=g({\mathbf{z}}_{t}).$ (1)
Note that ${\mathbf{z}}_{t}$ are divided into two parts, i.e., ${\mathbf{z}}_{t}=\{{\mathbf{z}}^{s}_{t},{\mathbf{z}}^{e}_{t}\}$, where ${\mathbf{z}}^{s}_{t}\in\mathbb{R}^{n_{s}}$ denote the environment-irrelated stationary latent variables, ${\mathbf{z}}^{e}_{t}\in\mathbb{R}^{n_{e}}$ denote the environment-related nonstationary latent variables, and $n_{e}+n_{s}=n$. Specifically, the $i$-th dimension stationary latent variable ${\mathbf{z}}_{t,i}^{s}$ are time-delayed and causally related to the historical stationary latent variables ${\mathbf{z}}_{t-\tau}^{s}$ with the time lag of $\tau$ via a nonparametric function $f_{i}^{s}$, which is formalized as follows:
$\small z_{t,i}^{s}\!=\!f^{s}_{i}(\{z^{s}_{t-\tau,k}|z^{s}_{t-\tau,k}\!\in\!\mathbf{Pa}(z_{t,i}^{s})\},\epsilon^{s}_{t,i})\ \ \text{with}\ \ \epsilon^{s}_{t,i}\sim p_{\epsilon_{i}^{s}},$ (2)
where $\mathbf{Pa}(z_{t,i}^{s})$ denotes the set of latent variables that directly cause ${\mathbf{z}}^{s}_{t,i}$ and $\epsilon^{s}_{t,i}$ denotes the temporally and spatially independent noise extracted from a distribution $p_{\epsilon_{i}^{s}}$. Moreover, the nonstationary latent variables ${\mathbf{z}}^{e}_{t}$ are influenced by the latent and discrete environment variables ${\mathbf{e}}_{t}$, which follows a first-order Markov process with $\bm{\mathrm{E}}\times\bm{\mathrm{E}}$ transition matrix $\bm{\mathrm{A}}$ and $\bm{\mathrm{E}}$ is the cardinality of ${\mathbf{e}}_{t}$. More specifically, we let the $(k,l)$-th entry $\bm{\mathrm{A}}_{k,l}$ be the probability from the state $k$ to the state $l$. As a result, the generation process of the $j$-th dimension nonstationary latent variable ${\mathbf{z}}_{t,j}^{e}$ can be formalized as follows:
$\vspace{-1.5mm}\small\left\{\begin{array}[]{l}{\mathbf{e}}_{1},{\mathbf{e}}_{2},\cdots,{\mathbf{e}}_{T}\sim\text{Markov Chain}(\mathbf{A})\\ z_{t,j}^{e}=f^{e}_{j}({\mathbf{e}}_{t},\epsilon^{e}_{t,j})\ \ \text{with}\ \ \epsilon^{e}_{t,j}\sim p_{\epsilon_{j}^{e}},\end{array}\right.\vspace{0.5mm}$ (3)
in which $f_{j}^{e}$ is a bijection function and $\epsilon^{e}_{t,j}$ is the mutually-independent noise extracted from $p_{\epsilon_{j}^{e}}$.
@@
To better understand the data generation process as shown in Figure 2, we provide an easily comprehensible example of human driving. First, we let ${\mathbf{x}}_{t}$ be the speed of a car. Then ${\mathbf{e}}_{t}$ denotes the action of the driver, i.e., speeding or braking, and ${\mathbf{z}}^{e}_{t}$ denotes the engine power or acceleration. Finally, ${\mathbf{z}}_{t}^{s}$ denotes road conditions such as flat and slippery roads, which are irrelated to the actions.
@@ 2.2 Identifying Distribution of Time Series Data for Nonstationary Time Series Forecasting
Based on the aforementioned data generation process, we aim to address the nonstationary time series forecasting problem, i.e., to predict the future observation $\{{\mathbf{x}}_{t+1},{\mathbf{x}}_{t+2},\cdots,{\mathbf{x}}_{T}\}$ only from the historical observation data $\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{t}\}$. Mathematically, our goal is to identify the joint distribution of the historical and future time series data. By combining the generation mechanism of Fig. 2, the joint distribution can be further derivated as follows:
$\vspace{-2mm}\small\begin{split}p({\mathbf{x}})=&\sum_{{\mathbf{e}}}\int_{{\mathbf{z}}^{e}}\int_{{\mathbf{z}}^{s}}p({\mathbf{x}},{\mathbf{e}},{\mathbf{z}}^{e},{\mathbf{z}}^{s})d_{{\mathbf{z}}^{e}}d_{{\mathbf{z}}^{s}}\\ =&\sum_{{\mathbf{e}}}\int_{{\mathbf{z}}^{e}}\int_{{\mathbf{z}}^{s}}p({\mathbf{x}}|{\mathbf{z}}^{e},{\mathbf{z}}^{s})p({\mathbf{z}}^{e}|{\mathbf{e}})p({\mathbf{e}})p({\mathbf{z}}^{s})d_{{\mathbf{z}}^{e}}d_{{\mathbf{z}}^{s}},\end{split}$ (4)
@@
where ${\mathbf{e}}:=\{{\mathbf{e}}_{1},\cdots,{\mathbf{e}}_{T}\},{\mathbf{z}}^{e}:=\{{\mathbf{z}}^{e}_{1},\cdots,{\mathbf{z}}^{e}_{T}\}$, and ${\mathbf{z}}^{s}:=\{{\mathbf{z}}^{s}_{1},\cdots,{\mathbf{z}}^{s}_{T}\}$ (we omit the subscripts due to limited space). Therefore, the joint distribution is identifiable by modeling the following four distributions: 1) the generative model of observations given stationary and nonstationary latent variables, i.e., $p({\mathbf{x}}|{\mathbf{z}}^{e},{\mathbf{z}}^{s})$; 2) the marginal distribution of latent environment variables, i.e. $p({\mathbf{e}})$ (Theorem 3.1); 3) the distribution of stationary latent variables,i.e., $p({\mathbf{z}}^{s})$ (Theorem 3.2); and 4) the conditional distribution of nonstationary latent variables, i.e., $p({\mathbf{z}}^{e}|{\mathbf{e}})$ (Theorem 3.3).
3 Identification of Latent Variables Under Sufficient Observation
In this section, we show the identifiability of these latent variables. To well establish the identification results of latent variables, we first illustrate the intuition behind the theories. Without any further assumptions, ${\mathbf{z}}_{t}^{s}$ and ${\mathbf{z}}_{t}^{e}$ are symmetric with ${\mathbf{x}}_{t}$, implying the simultaneous changing of ${\mathbf{z}}_{t}^{s}$ and ${\mathbf{z}}_{t}^{e}$, which makes it hard to disentangle ${\mathbf{z}}_{t}^{s}$ and ${\mathbf{z}}_{t}^{e}$ since ${\mathbf{z}}_{t}^{e}$ is mutually dependent without given ${\mathbf{e}}_{t}$. To break this restriction, we propose the Sufficient Observations assumption, which requires at least 2 consecutive observations for each environment. Therefore, for any three consecutive observations ${\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3}$, the case where their corresponding environments are different, that is, ${\mathbf{e}}_{1}\neq{\mathbf{e}}_{2}\neq{\mathbf{e}}_{3}$, can be excluded. As a result, we can disentangle ${\mathbf{z}}_{t}^{s}$ and ${\mathbf{z}}_{t}^{e}$. In the following, we first show how to identify latent environments with Theorem 3.1, which addresses when the temporal distribution shift occurs. Then we show how to disentangle the stationary and nonstationary latent variables, i.e., ${\mathbf{z}}_{t}^{s}$ and ${\mathbf{z}}_{t}^{e}$, with Theorem 3.2 and 3.3, which addresses how the stationary and nonstationary latent variables change over time. Please refer to Appendix B for more details on the definition of identification up to label swapping and component-wise identification as well as the subspace identification.
3.1 Identification of Latent Environment Variables ${\mathbf{e}}_{t}$
To identify when the temporal distribution shift occurs and estimate the marginal distribution $p({\mathbf{e}})$, we propose the identification theory of latent environment variables ${\mathbf{e}}_{t}$, which is shown as Theorem 3.1.
@@ 3.1 Identification of Latent Environment Variables ${\mathbf{e}}_{t}$
(Identification of the latent environment ${\mathbf{e}}_{t}$.) Suppose the observed data is generated following the data generation process in Figure 2 and Equation (1)-(3). Then we further make the following assumptions:
  • A3.1.1 (Prior Environment Number:) The number of latent environments of the Markov process, $\bm{\mathrm{E}}$, is known.
  • A3.1.2 (Full Rank:) The transition matrix $\mathbf{A}$ is full rank.
  • A3.1.3 (Linear Independence:) For ${\mathbf{e}}_{t}=1,2,\cdots,E$, the probability measures $\mu_{e}=p({\mathbf{x}}_{t}|{\mathbf{z}}_{t}^{e},{\mathbf{e}}_{t})$ are linearly independence and for any two different probability measures $\mu_{i},\mu_{j}$, their ratio $\frac{\mu_{i}}{\mu_{j}}$ are linearly independence.
  • A3.1.4 (Sufficient Observation:) For each ${\mathbf{e}}_{t}=\{1,\!\cdots,\!\bm{\mathrm{E}}\}$, there are at least 2 consecutive observations.
Proof Sketch. First, given any three consecutive observations ${\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3}$ with the corresponding latent environments ${\mathbf{e}}_{1},{\mathbf{e}}_{2},{\mathbf{e}}_{3}$, there are only three cases of the latent environment variables with the help of the sufficient observations, and they are 1) ${\mathbf{e}}_{1}={\mathbf{e}}_{2}={\mathbf{e}}_{3}$ 2) ${\mathbf{e}}_{1}={\mathbf{e}}_{2}\neq{\mathbf{e}}_{3}$ 3) ${\mathbf{e}}_{1}\neq{\mathbf{e}}_{2}={\mathbf{e}}_{3}$, and we discuss these cases, respectively. For each case, we derive the joint distribution of $p({\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3})$ to the product of three independent measures w.r.t. $p({\mathbf{e}}_{2})$. Sequentially, by employing the extension of Kruskal’s theorem (Kruskal, 1977, 1976), i.e., the Theorem 9 of (Allman et al., 2009), the latent environment variables can be detected with identification guarantees. The detailed proof of Theorem 3.1 is provided in Appendix E.1.
Discussion of the Assumptions. To show the reasonableness of our theoretical result, we further discuss the intuition of the assumptions. First, the prior environment number assumption means that we can take the number of environments as prior knowledge, for example, we can know the number of actions of drivers. Then the full rank assumption implies that any environment is accessible from other environments. The Linear Independent assumption shows that the influence from each ${\mathbf{e}}_{t}$ to ${\mathbf{x}}_{t}$ is sufficiently different, for example, how speeding and braking influence the speed of a car is totally different. Finally, the sufficient observations assumption is easy to meet since each action should last for an extended duration, and we can achieve it with a sufficiently high sampling frequency in real-world scenarios.
3.2 Identification of Stationary Latent Variables ${\mathbf{z}}^{s}_{t}$
Then we prove that the stationary latent variables ${\mathbf{z}}^{s}_{t}$ are subspace identifiable with the help of nonlinear ICA.
Theorem 3.2.
(Identification of the stationary latent variables ${\mathbf{z}}^{s}_{t}$.) We follow the data generation process in Figure 2 and Equation (1)-(3) and make the following assumptions:
  • A3.2.1 ((Smooth and Positive Density:)) The probability density function of latent variables is smooth and positive, i.e. $p({\mathbf{z}}_{t}^{s}|{\mathbf{z}}_{t-1}^{s})>0$ over $\mathcal{Z}_{t}$ and $\mathcal{Z}_{t-1}$.
  • A3.2.2 (Conditional independent:) Conditioned on ${\mathbf{z}}_{t-1}^{s}$, each ${\mathbf{z}}_{t,i}^{s}$ is independent of any other ${\mathbf{z}}_{t,j}^{s}$ for $i,j\in\{n_{e}+1,\cdots,n\},i\neq j$, i.e., $\log p({\mathbf{z}}_{t}^{s}|{\mathbf{z}}^{s}_{t-1})=\sum_{i=n_{e}+1}^{n}\log p(z_{t,i}^{s}|{\mathbf{z}}_{t-1}^{s})$.
  • A3.2.3 (Linear independence): For any ${\mathbf{z}}_{t}^{s}\in\mathcal{Z}^{s}\subseteq\mathbb{R}^{n_{s}}$, there exist $n_{s}+1$ values of $z^{s}_{t-1,l},l=n_{e}+1,\cdots,n$, such that these $n_{s}$ vectors $\mathbf{v}_{t,k,l}-\mathbf{v}_{t,k,n}$ are linearly independent, where $\mathbf{v}_{t,k,l}$ is defined as follows:
    $\small\mathbf{v}_{t,k,l}=\left(\frac{\partial\log p(z_{t,k}^{s}|{\mathbf{z}}^{s}_{t-1})}{\partial z_{t,k}^{s}\partial z_{t-1,n_{e}+1}^{s}},\cdots,\frac{\partial\log p(z_{t,k}^{s}|{\mathbf{z}}^{s}_{t-1})}{\partial z_{t,k}^{s}\partial z_{t-1,n-1}^{s}}\right)$ (5)
Then, by learning, the data generation process ${\mathbf{z}}^{s}_{t}$ is subspace identifiable.
Proof Sketch. First, we construct an invertible transformation $h$ between the ground-truth latent variables ${\mathbf{z}}_{t}$ and estimated $\hat{{\mathbf{z}}}_{t}$. Then we employ the variance of different environments to construct a full-rank linear system, where the only solution of $\frac{\partial{\mathbf{z}}^{e}_{t}}{\partial\hat{{\mathbf{z}}}^{s}_{t}}$ is zero. Because the Jacobian of $h$ is invertible, for each $z_{t,i}^{s},i\in\{n_{e}+1,\cdots,n\}$, there exists a $h_{i}$ such that $z^{s}_{t,i}=h_{i}(\hat{{\mathbf{z}}}^{s}_{t})$ and ${\mathbf{z}}^{s}_{t}$ is subspace identifiable.
Discussion of the Assumptions. The proof can be found in the Appendix E.2. The first two assumptions are standard in the component-wise identification of existing nonlinear ICA (Kong et al., 2022; Khemakhem et al., 2020a). The third assumption implies that $p(z^{s}_{t,k}|{\mathbf{z}}^{s}_{t-1})$ should change sufficiently over ${\mathbf{z}}^{s}_{t-1}$, which can be easily satisfied by sampling sufficiently. Compared with existing component-wise identification results in time series modeling (Yao et al., 2021, 2022; Song et al., 2023), which requires $12n_{s}+$ 1 auxiliary variables, our subspace identification result yields a similar result but with a milder condition.
3.3 Identification of Nonstationary Latent Variables ${\mathbf{z}}^{e}_{t}$
Theorem 3.3.
@@ 3.3 Identification of Nonstationary Latent Variables ${\mathbf{z}}^{e}_{t}$
(Identification of the nonstationary latent variables ${\mathbf{z}}^{e}_{t}$.) We follow the data generation process in Figure 2 and Equation (1)-(3), then we make the following assumptions:

### v4

# Nonstationary Time Series Forecasting via Unknown Distribution Adaptation

@@ Abstract
As environments evolve, temporal distribution shifts can degrade time series forecasting performance. A straightforward solution is to adapt to nonstationary changes while preserving stationary dependencies. Hence some methods disentangle stationary and nonstationary components by assuming uniform distribution shifts, but it is impractical since when the distribution changes is unknown. To address this challenge, we propose the Unknown Distribution Adaptation (UDA) model for nonstationary time series forecasting, which detects when distribution shifts occur and disentangles stationary/nonstationary latent variables, thus enabling adaptation to unknown distribution without assuming a uniform distribution shift. Specifically, under a Hidden Markov assumption of latent environments, we demonstrate that the latent environments are identifiable. Sequentially, we further disentangle stationary/nonstationary latent variables by leveraging the variability of historical information. Based on these theoretical results, we propose a variational autoencoder-based model, which incorporates an autoregressive hidden Markov model to estimate latent environments. Additionally, we further devise the modular prior networks to disentangle stationary/nonstationary latent variables. These two modules realize automatic adaptation and enhance nonstationary forecasting performance. Experimental results on several datasets validate the effectiveness of our approach.
@@ 1 Introduction
Time series forecasting (Zhou et al.,, 2021; Lim and Zohren,, 2021; Rangapuram et al.,, 2018; Chatfield,, 2000; Zhang,, 2003) has achieved pioneering applications in various fields (Bi et al.,, 2023; Wu et al.,, 2023; Sezer et al.,, 2020). However, the inherent nonstationarity of time series data hinders the forecasting models from generalizing on the temporally varying distribution shift.
Several methodologies are proposed to solve this problem, which can be categorized into two types according to the inter-instance and intra-instance temporal distribution shift assumptions. The first type of method assumes that the shift occurs among instances, and each sequence instance is stationary (Li et al., 2023b, ; Oreshkin et al.,, 2021). Therefore, instance normalization (Kim et al.,, 2021) or nonstationary attention mechanism (Liu et al.,, 2022) is used to remove nonstationary components and compensate for them in prediction. Another type assumes that the environment changes uniformly (Liu et al., 2023c, ; Surana,, 2020). Therefore, some researchers adopt stationarization (Virili and Freisleben,, 2000) to remove nonstationarity from time series data. And (Liu et al.,, 2022) partition the time-series data into equally-sized and stationary segments and uses the Fast Fourier Transform to select stationary and nonstationary components. Recent advances (Liu et al., 2024a, ) further learn the invariant information to achieve out-of-distribution generalization, but they neglect the environment-related information. In summary, these methods aim to disentangle the stationary and nonstationary dependency. More discussion about related works can be found in Appendix A.
Although these methods mitigate the temporal distribution shift to some extent, the assumptions they require are usually too strict since each time series instance or segment may not be stationary, especially when latent environment changes are unknown. Figure 1 illustrates an example where a nonstationary sine curve is influenced by the nonstationary latent variable (amplitude) and stationary latent variables (frequency and phase). Just like the example in Figure 1 (a), existing methods assume a uniform distribution shift and partition the nonstationary time series into three equal segments. However, the purple and green curves remain nonstationary, making disentangling stationary and nonstationary latent variables challenging. Besides correct latent environment estimations, proper disentanglement is crucial for adapting to distribution changes. Figure 1 (b) shows that if latent variables are entangled—such as mixing amplitude and phase—the model struggles to preserve the stationary dependencies and update the nonstationary ones.
@@
The above points emphasize the importance of a nonstationary time series forecasting model capable of accurately identifying environment changes and disentangling stationary and nonstationary latent variables, as illustrated in Figure 1 (c). Standing on this insight, we therefore establish the Unknown Distribution Adaptation model (UDA in short), which offers identification guarantees to ensure the performance of nonstationary time series forecasting. Specifically, under the assumption that latent environments follow a Hidden Markov process, we first prove that these environments can be identified from observations. Sequentially, by harnessing the variability of historical information, we can further identify the stationary and nonstationary latent variables, ensuring the proper adaptation to new environments. Guided the theoretical results, we develop the UDA model based on variational autoencoder (VAE). The UDA model is equipped with an autoregressive hidden Markov model for environment estimation and modular prior networks for stationarity/nonstationarity disentanglement. Evaluation of simulation and eight real-world benchmark datasets demonstrates the accuracy of latent environment estimation and identification of latent states, as well as the effectiveness of real-world applications.
2 Problem Setup
@@ 2.1 Data Generation Process for Time Series Data
To illustrate how we address the nonstationary time series forecasting problem, we begin with the data generation process as shown in Figure 2. Suppose that we have time series data with discrete time steps, $\textbf{X}=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{T}\}$, where ${\mathbf{x}}_{t}\in\mathbb{R}$ are generated from latent variables ${\mathbf{z}}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{n}$ by an invertible and non-linear mixing function $g$ as shown in Equation (1)
${\mathbf{x}}_{t}=g({\mathbf{z}}_{t}).$ (1)
Note that ${\mathbf{z}}_{t}$ are divided into two parts, i.e., ${\mathbf{z}}_{t}=\{{\mathbf{z}}^{s}_{t},{\mathbf{z}}^{e}_{t}\}$, where ${\mathbf{z}}^{s}_{t}\in\mathbb{R}^{n_{s}}$ denote the environment-irrelated stationary latent variables, ${\mathbf{z}}^{e}_{t}\in\mathbb{R}^{n_{e}}$ denote the environment-related nonstationary latent variables, and $n_{e}+n_{s}=n$. Specifically, the $i$-th dimension stationary latent variable $z_{t,i}^{s}$ is time-delayed and causally related to the historical stationary latent variables ${\mathbf{z}}_{t-\tau}^{s}$ with the time lag of $\tau$ via a nonparametric function $f_{i}^{s}$, which is formalized as follows:
$z_{t,i}^{s}\!=\!f^{s}_{i}(\{z^{s}_{t-\tau,k}|z^{s}_{t-\tau,k}\!\in\!\mathbf{Pa}(z_{t,i}^{s})\},\varepsilon^{s}_{t,i})\ \ \text{with}\ \ \varepsilon^{s}_{t,i}\sim p_{\varepsilon_{i}^{s}},$ (2)
where $\mathbf{Pa}(z_{t,i}^{s})$ denotes the set of latent variables that directly cause ${\mathbf{z}}^{s}_{t,i}$ and $\varepsilon^{s}_{t,i}$ denotes the temporally and spatially independent noise extracted from a distribution $p_{\varepsilon_{i}^{s}}$. Moreover, the nonstationary latent variables ${\mathbf{z}}^{e}_{t}$ are influenced by the latent and discrete environment variables ${\mathbf{e}}_{t}$, which follow a first-order Markov process with $\bm{\mathrm{E}}\times\bm{\mathrm{E}}$ transition matrix $\bm{\mathrm{A}}$ and $\bm{\mathrm{E}}$ is the cardinality of ${\mathbf{e}}_{t}$. More specifically, we let the $(k,l)$-th entry $\bm{\mathrm{A}}_{k,l}$ be the probability from the state $k$ to the state $l$. As a result, the generation process of the $j$-th dimension nonstationary latent variable ${\mathbf{z}}_{t,j}^{e}$ can be formalized as:
$\left\{\begin{array}[]{l}{\mathbf{e}}_{1},{\mathbf{e}}_{2},\cdots,{\mathbf{e}}_{T}\sim\text{Markov Chain}(\mathbf{A})\\ z_{t,j}^{e}=f^{e}_{j}({\mathbf{e}}_{t},\varepsilon^{e}_{t,j})\ \ \text{with}\ \ \varepsilon^{e}_{t,j}\sim p_{\varepsilon_{j}^{e}},\end{array}\right.$ (3)
in which $f_{j}^{e}$ is a bijection function and $\varepsilon^{e}_{t,j}$ is the mutually-independent noise extracted from $p_{\varepsilon_{j}^{e}}$.
@@
To better understand the data generation process in Figure 2, we provide a comprehensible example of human driving. First, we let ${\mathbf{x}}_{t}$ be the speed of a car. Then ${\mathbf{e}}_{t}$ denotes the action of the driver, i.e., speeding or braking, and ${\mathbf{z}}^{e}_{t}$ denotes the engine power or acceleration. Finally, ${\mathbf{z}}_{t}^{s}$ denotes road conditions such as flat and slippery roads, which are irrelated to the actions.
@@ 2.2 Identifying Distribution of Time Series Data for Nonstationary Time Series Forecasting
Based on this data generation process, we aim to address the nonstationary time series forecasting problem, i.e., to predict the future observation $\{{\mathbf{x}}_{t+1},{\mathbf{x}}_{t+2},\cdots,{\mathbf{x}}_{T}\}$ only from the historical observation data $\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{t}\}$. Mathematically, our goal is to identify the joint distribution of the historical and future time series data. By combining the generation mechanism of Fig. 2, the joint distribution can be further derived as follows:
$\begin{split}p(\textbf{X})&=\sum_{{\mathbf{e}}}\int_{{\mathbf{z}}^{e}}\int_{{\mathbf{z}}^{s}}p(\textbf{X},{\mathbf{e}},{\mathbf{z}}^{e},{\mathbf{z}}^{s})d_{{\mathbf{z}}^{e}}d_{{\mathbf{z}}^{s}}\\ &=\sum_{{\mathbf{e}}}\int_{{\mathbf{z}}^{e}}\int_{{\mathbf{z}}^{s}}p(\textbf{X}|{\mathbf{z}}^{e},{\mathbf{z}}^{s})p({\mathbf{z}}^{e}|{\mathbf{e}})p({\mathbf{e}})p({\mathbf{z}}^{s})d_{{\mathbf{z}}^{e}}d_{{\mathbf{z}}^{s}},\end{split}\vspace{-3mm}$ (4)
@@
where ${\mathbf{e}}:=\{{\mathbf{e}}_{1},\cdots,{\mathbf{e}}_{T}\},{\mathbf{z}}^{e}:=\{{\mathbf{z}}^{e}_{1},\cdots,{\mathbf{z}}^{e}_{T}\}$, and ${\mathbf{z}}^{s}:=\{{\mathbf{z}}^{s}_{1},\cdots,{\mathbf{z}}^{s}_{T}\}$ (we omit the subscripts due to limited space). Therefore, the joint distribution is determined by modeling the following four distributions: 1) the generative model of observations given stationary and nonstationary latent variables, i.e., $p({\mathbf{x}}|{\mathbf{z}}^{e},{\mathbf{z}}^{s})$; 2) the marginal distribution of latent environment variables, i.e., $p({\mathbf{e}})$; 3) the distribution of stationary latent variables,i.e., $p({\mathbf{z}}^{s})$; and 4) the conditional distribution of nonstationary latent variables, i.e., $p({\mathbf{z}}^{e}|{\mathbf{e}})$.
3 Identification of Latent Variables
In this section, we show the identifiability  [^1] Please refere to Appendix B for the definition of different types of identification. of these latent variables. To well establish the identification results of latent variables, we first leverage Theorem 3.1 to show that the stationary and nonstationary latent variables are block-wise identifiable by employing the fact that ${\mathbf{z}}^{e}_{t+1}$ depend on ${\mathbf{z}}_{t-1}^{e}$ given ${\mathbf{x}}_{t}$. Sequentially, we show that the transition of the latent environments is identifiable up to label swapping, which is shown in Lemma 3.2. Finally, we prove that stationary and nonstationary latent variables are component-wise identifiable, which is shown in the Lemma 3.3.
3.1 Latent Environment Identification
To partition the stationary latent variables ${\mathbf{z}}_{t}^{s}$ and nonstationary latent variables ${\mathbf{z}}_{t}^{e}$, we propose the block-wise identification theory, which is shown in Theorem 3.1.
@@ 3.1 Latent Environment Identification
(Block-wise identifiability of the nonstationary latent variables ${\mathbf{z}}_{t}^{e}$ and the stationary latent variables ${\mathbf{z}}_{t}^{s}$.) We follow the data generation process in Figure 2 and Equation (1)-(3), then we make the following assumptions: